Question 2: Supply Suppose that Jessica faces the following total cost function: C(y) = 2y2 + y + 200 Note that 200 denotes t(b) Derive the short run supply function for Jessica. Suppose p = 61. How much y does Jessica supply in the short run? What i(c) Derive the long run supply function for Jessica. Suppose p= 61. How much output does Jessica supply in the long run? What(d) Both in short run and long run, Jessica supplies a strictly positive amount of y(i.e. y > 0). True or False? Explain why.
Answer #1

Poee Groen Total Cost function CCY) = 2y?+4200 fixed cost 200 ca Marginal cost MC=Hyt Average cost <A C= 2y +1+200ly Averagesetz | 200 | 8 = as م 14 = 1 + (38 aols+ 1+ (38 3= أ+ (8) قہر - 16 = 1 + ( aca ) + + 2ools aq 37: 1 + (@ با 24 - 2009 ه aootbCost 1 2 3 н5618 0 10 11 12 13 14 Б % ---MC ---- AC - AVC- * AFC оо1 P (9)Page 3 b) F-61, out put Soppted ewals Асар u3+1 = 61 uy=60 19=15 Jessica Pooducer su oplos - total revenue - total valable coPage the Jessica porofit = to fal gevenue -tolal cost = 61X10 -2 x 100+10+200 = 610-410 Jessica Pootit = 200) d) yes, Jeffca